OpenStudy (falloutboy):

Find the missing measure for a right circular cone given the following information. Find L.A. if r = 6 and h = 8 a048 b)60 c)96

OpenStudy (toriacm18):

volume = 1/3 * pi *r^2*h = 96pi plug in all your known values and solve for the missing variable

OpenStudy (toriacm18):

c

OpenStudy (pawanyadav):

Here slant height l^2= 64+36=100 So l=√100=10